How far from a piece of paper must you hold your fathers +2D reading glasses to try to burn a hole in the paper with sunlight?
Answer #1

+2D=1/f

f=1/2 metre=50cm

Paper should be kept at 50 cm

(rays coming from infinity are converged at focus)